# Iterative Prompt Relocation for Distribution-Adaptive Visual Prompt Tuning

### v1

@@ Abstract
Visual prompt tuning (VPT) provides an efficient and effective solution for adapting pre-trained models to various downstream tasks by incorporating learnable prompts. However, most prior art indiscriminately applies a fixed prompt distribution across different tasks, neglecting the importance of each block differing depending on the task. In this paper, we investigate adaptive distribution optimization (ADO) by addressing two key questions: (1) How to appropriately and formally define ADO, and (2) How to design an adaptive distribution strategy guided by this definition? Through in-depth analysis, we provide an affirmative answer that properly adjusting the distribution significantly improves VPT performance, and further uncover a key insight that a nested relationship exists between ADO and VPT. Based on these findings, we propose a new VPT framework, termed PRO-VPT (iterative Prompt RelOcation-based VPT), which adaptively adjusts the distribution building upon a nested optimization formulation. Specifically, we develop a prompt relocation strategy for ADO derived from this formulation, comprising two optimization steps: identifying and pruning idle prompts, followed by determining the optimal blocks for their relocation. By iteratively performing prompt relocation and VPT, our proposal adaptively learns the optimal prompt distribution, thereby unlocking the full potential of VPT. Extensive experiments demonstrate that our proposal significantly outperforms state-of-the-art VPT methods, e.g., PRO-VPT surpasses VPT by 1.6% average accuracy, leading prompt-based methods to state-of-the-art performance on the VTAB-1k benchmark. The code is available at https://github.com/ckshang/PRO-VPT.
@@ 1 Introduction
Fine-tuning pre-trained vision models (PVMs) [24] has proven remarkably effective in adapting to a variety of downstream tasks [58, 14]. However, the computational and storage costs associated with full fine-tuning are prohibitively high, particularly as model sizes continue to grow [37, 52, 1]. To overcome these challenges, parameter-efficient fine-tuning (PEFT) methods [55, 56, 18, 2], such as visual prompt tuning (VPT) [20], have emerged as more promising alternatives, garnering significant attention.
@@
Instead of updating all model parameters directly, VPT inserts learnable prompt tokens into the input space of Transformer blocks, with either a shallow distribution (inserted only in the first block) or a deep distribution (uniformly inserted across all blocks), while keeping the PVM frozen. This approach significantly reduces the number of parameters that need to be fine-tuned, thereby lowering the computational and memory overhead [14]. Follow-up research further enhances this pipeline by improving aspects such as prompting architecture [13, 46], prompt initialization [53], and the propagation structures of prompts [62, 57, 5].
@@
Despite these advancements, most of the prior art still relies on a preset fixed prompt distribution, either shallow or deep, across various downstream tasks, as illustrated in Fig. 1. However, recent findings reveal that the importance of each block varies significantly across different tasks [45, 41, 12, 26, 54], underscoring the need to dynamically tailor the fine-tuning intensity for each block to improve task performance [27, 48, 11, 42, 23]. This implies that indiscriminately using a fixed prompt distribution is inadequate, motivating us to adaptively adjust the distribution to better accommodate each task.
@@
In this paper, we naturally consider prompt distribution as one of the optimization objectives, which we refer to as adaptive distribution optimization (ADO). As the focus pivots towards the ADO problem, new questions arise. First, given the unexplored nature of prompt distribution optimization, the definition of ADO remains unclear. For instance, defining it as a one-shot process or an iterative one leads to a critical consideration, namely whether the effectiveness of distribution adjustments is influenced by ongoing prompt updates. This leads us to our first key question: How can we appropriately and formally define ADO? Second, once the definition of ADO is established, a more essential question emerges: How can we achieve adaptive distribution derived from this definition?
@@
To address these questions, we first empirically analyze the ADO problem. Our findings confirm the significance of ADO and further highlight two key insights: the effectiveness of distribution adjustments is indeed impacted by prompt updates, and a nested relationship exists between ADO and VPT. To this end, we propose a novel iterative framework, termed PRO-VPT (iterative Prompt RelOcation-based VPT), with a nested optimization definition for ADO and VPT. Specifically, building on this definition, we develop a prompt relocation (PR) strategy for adaptively refining the distribution. It comprises two key optimization steps: pruning and allocating. The pruning step first identifies and releases idle prompts, while the subsequent allocation step determines optimal blocks for redistributing these prompts. By iteratively executing PR and VPT, the proposed PRO-VPT adaptively learns the optimal prompt distribution for each task, unlocking the full potential of VPT and leading to improved robustness. Our contributions can be summarized as follows:
@@
  • Derived from the definition, we present a PR strategy for ADO, which introduces an idleness score to prune underutilized prompts and allocate them to more effective blocks using reinforcement learning (RL).
  • Extensive experiments show that PRO-VPT reaches the state-of-the-art average accuracy of 78.0% on VTAB-1k, with a 1.6% advantage over the VPT baseline.
@@ 2 Related Work
Visual Prompt Tuning. VPT [20], one of the leading techniques in PEFT, draws inspiration from language-domain prompting works [37, 34, 29, 35] to fine-tune a small portion of prompts while keeping the pre-trained vision Transformer (ViT) frozen. Subsequent research has improved VPT in three main aspects: (1) Extending prompts to other Transformer components (e.g., key and value matrices or the patch embedding layer) to better suit the Transformer architecture [13, 46]; (2) Initializing prompts with prototypes of image tokens to accelerate convergence [53]; (3) Refining the propagation structures of prompts to strengthen cross-layer interactions [62, 57, 5]. However, most existing methods depend on a fixed prompt distribution across different tasks, resulting in less effective prompting solutions. In contrast, our approach adaptively adjusts the prompt distribution for each task, enabling prompt-based methods to achieve advanced performance.
Layer-wise fine-tuning. Resent research [49, 28, 12] has highlighted the importance of recognizing the varying contributions of pre-trained blocks for enhancing fine-tuning performance and robustness. For instance, [27] demonstrates that identifying pre-trained blocks that are already near-optimal for the downstream task and selectively fine-tuning the remaining blocks can significantly improve robustness to distribution shifts. Similarly, [45] shows that assigning different learning rates for each block is advantageous for few-shot learning. Furthermore, TPGM [48] introduces a bi-level optimization to optimize a projection constraint for each block, aiming to learn how much to fine-tune each block. Inspired by these findings, we hypothesize that the prompting intensity for each block, i.e., the overall distribution of prompts, may also play a crucial role in task adaptation, leading us to focus on adaptively adjusting the prompt distribution within VPT.
@@ 3 Methodology
3.1 Motivation
Many previous studies have emphasized the significance of treating each Transformer block differently during downstream task fine-tuning [45, 27, 12, 23]. For instance, a notable study in [27] reveals that certain blocks pre-trained on the upstream task may already be near-optimal for specific downstream tasks, hence over-tuning could impair generalization. As a result, recent advances have focused on developing adaptive fine-tuning strategies to calibrate the fine-tuning intensity for each block, which have shown promising potential in improving both performance and robustness [41, 49]. This suggests that utilizing a fixed prompt distribution (essentially the number of prompts/prompting intensity per block) for all tasks, without any differentiation, might not fully exploit the potential of VPT. Motivated by this insight, our work focuses on optimizing the prompt distribution to better accommodate downstream tasks and enhance overall robustness (cf. § 4.5).
@@ 3.2 Analysis and Formulation of ADO
We first examine the underlying nature of ADO and its precise formulation. Intuitively, ADO can be defined in two distinct ways: one-shot (independently adjusting the distribution) or iterative (integrating with prompt tuning through iteratively adjusting the distribution). These definitions are distinct in whether the distribution is adjusted concurrently with prompt tuning. Specifically, an iterative process is more suitable if distribution adjustments are influenced by ongoing prompt updates, as independent optimization might lead to suboptimal results without accounting for these changes. Conversely, one-shot adjustments may suffice if the adjustments remain unaffected by prompt updates.
@@
To investigate which definition better matches the underlying nature, we experiment to observe how distribution adjustments perform when applied to prompts from different epochs. We use two intuitive adjustment strategies: incrementally adding to a small initial prompt set and repositioning existing prompts. As shown in Fig. 2, adding or repositioning prompts to specific blocks can effectively improve performance; however, the effective adjustments differ significantly when prompts are updated. Comprehensive analysis with detailed results is provided in Appendix B. In conclusion, we summarize our key discoveries as follows:
Discovery 1. Properly adjusting the prompt distribution significantly enhances performance, confirming the necessity to reconsider its optimization.
Discovery 2. The effectiveness of distribution adjustments is sensitive not only to different blocks but also to the updated prompts themselves, highlighting the need for an iterative process that continuously adjusts the distribution over time.
Discovery 3. Distribution adjustments can only be properly evaluated after prompt tuning, indicating that the prompt tuning process should be nested within the distribution optimization process (see Appendix B).
Building on these insights, we propose integrating ADO with VPT into an iterative optimization framework. Formally, we present the definition of this ADO-VPT co-design framework through the lens of nested optimization. Let $\mathbf{P}\!=\!\{\mathbf{p}_{k}\}^{N}_{k=1}$ denote a set of $N$ prompts and $\mathcal{D}\!=\!\{d_{k}\}^{N}_{k=1}$ represent their distributed blocks, the objectives of ADO and VPT can then be expressed as:
$\mathcal{D}^{\ast}$ $=\mathop{\arg\min}\limits_{\mathcal{D}}\leavevmode\nobreak\ \mathbb{E}_{(\mathbf{x},y)\in\mathcal{T}_{tr}}\big{[}\mathcal{L}(f_{\mathbf{P}^{\ast},\mathcal{D}}(\mathbf{x}),y)\big{]},$ (1)
$\mathbf{P}^{\ast}$ $=\mathop{\arg\min}\limits_{\mathbf{P}}\leavevmode\nobreak\ \mathbb{E}_{(\mathbf{x},y)\in\mathcal{T}_{tr}}\big{[}\mathcal{L}(f_{\mathbf{P},\mathcal{D}^{\ast}}(\mathbf{x}),y)\big{]},$ (2)
where $d_{k}\!\in\!\{0\}\!\bigcup[L]$ ($d_{k}\!=\!0$ for undistributed) and $L$ denotes the number of Transformer blocks $\{B_{i}\}^{L}_{i=1}$. $\mathcal{T}_{tr}$ is the training set with input-label pairs $(\mathbf{x},y)$, $\mathcal{L}$ is the task loss function, and $f_{\mathbf{P},\mathcal{D}}$ is the prompted model. This formulation clearly illustrates the intertwined relationship between visual prompts $\mathbf{P}$ and their distribution $\mathcal{D}$, with its workflow presented in Fig. iii. A more detailed explanation for this nested optimization formulation is provided in Appendix B.
@@
To solve this nested optimization, we employ the alternating optimization approach [1], establishing an iterative process that alternates between ADO and VPT. Specific to ADO, we empirically observe that adding-based adjustments may lead to undertraining of newly introduced prompts and conflicts with existing ones, as illustrated in Appendix C. Therefore, we particularly focus on developing a prompt relocation (PR) strategy derived from the formulation to effectively tackle the ADO problem.
@@ 3.3 Iterative Prompt Relocation-based VPT
Since the objective of PR defined in Eq. (1) poses a discrete optimization problem that cannot be directly solved by backpropagation, we first introduce the PR strategy derived from the definition, followed by a detailed overview of the PRO-VPT framework. Building on our discoveries regarding ADO, the discrete PR objective is further complicated by the non-stationary nature of effective relocation (Discovery 2) and the necessity to predict the effectiveness of various relocation operations (Discovery 3). Nevertheless, these three properties make a RL framework [30, 22] a natural and well-suited approach for optimizing the PR objective, where all possible relocation operations are treated as actions.
However, a naive RL implementation considering all $L^{2}$ possible relocation arrangements (e.g., 144 actions for ViT-B) would require numerous iterations to achieve convergence (cf. § 4.3). To address this, we propose decomposing the PR process into two sequential steps: pruning and allocating. The pruning step first identifies and releases idle prompts for relocation, while the allocating step then employs RL to determine optimal blocks for these freed prompts. By focusing merely on possible allocatable blocks for the pruned prompts, this approach effectively reduces the action space from $L^{2}$ to $L$. Formalizing this, the optimization objectives for pruning and allocation, derived from Eq. (1), are:
$k^{\ast}=$ $\mathop{\arg\max}\limits_{k}\leavevmode\nobreak\ \mathbb{E}_{(\mathbf{x},y)\in\mathcal{T}_{tr}}\big{[}\Delta\mathcal{L}(f_{\mathbf{P},\mathcal{D}},f_{\mathbf{P},\mathcal{D}\mid d_{k}=0})\big{]},$ (3)
$a^{\ast}=$ $\mathop{\arg\max}\limits_{a}\leavevmode\nobreak\ \mathbb{E}_{(\mathbf{x},y)\in\mathcal{T}_{tr}}\big{[}\Delta\mathcal{L}(f_{\mathbf{P},\mathcal{D}^{-}},f_{\mathbf{P}^{\prime},\mathcal{D}\mid d_{k^{\!\ast\!}}=a})\big{]},$ (4)
where $\Delta\mathcal{L}(f_{1},f_{2})\!=\!\mathcal{L}(f_{1}(\mathbf{x}),y)\!-\!\mathcal{L}(f_{2}(\mathbf{x}),y)$ denotes the loss difference, $\mathcal{D}^{-}\!=\!\{d_{k}\!\in\!\mathcal{D}\!\mid\!d_{k^{\!\ast\!}}\!=\!0\}$ and $\mathcal{D}^{+}\!=\!\{d_{k}\!\in\!\mathcal{D}\!\mid\!d_{k^{\!\ast\!}}\!=\!a^{\ast}\}$ represent the prompt distributions after pruning and allocation, with $a\!\in\![L]$ as the allocation block index. In accordance with Discovery 3, Eq. (4) evaluates the allocation step using the tuned prompts $\mathbf{P}^{\prime}$. Remarkably, both pruning and allocation steps operate on a single prompt per iteration, with further explanation provided in Appendix D.
Next, we approach these two stages around the optimization objectives as follows.
Idleness-Score-based Pruning. To quantify whether a prompt is idle and suitable for relocation, we introduce an idleness score $\mathcal{I}_{k}$ based on Eq. (3):
@@
Clearly, if the idleness score $\mathcal{I}_{k}\!=\!\Delta\mathcal{L}\!>\!0$, the prompt $\mathbf{p}_{k}$ is more advantageous to prune than to retain, indicating that it is negative for the current model $f_{\mathbf{P},\mathcal{D}}$ and thereby can be pruned for potential relocation. Conversely, if $\mathcal{I}_{k}\!\leq\!0\leavevmode\nobreak\ \forall k$, no prompt is sufficiently idle to warrant removal.
Furthermore, since computing $\mathcal{I}_{k}$ for each individual prompt $\mathbf{p}_{k}$ is significantly inefficient, we approximate $\mathcal{I}_{k}$ using its first-order Taylor expansion (details are available in Appendix E):
^$\widehat{\mathcal{I}}_{k}\approx\mathbf{g}_{k}^{T}\mathbf{p}_{k},$ (6)
@@
Therefore, by calculating the estimated scores as in Eq. (6), we can effectively identify and release the idle prompt $\mathbf{p}_{k^{\!\ast\!}}$ if ^$\widehat{\mathcal{I}}_{k^{\!\ast\!}}\!=\!\max\widehat{\mathcal{I}}_{k}\!>\!0$, thereby forming $\mathcal{D}^{-}$.
RL-based Allocation. After pruning an idle prompt, the following key step is to leverage RL to determine its optimal allocation to the block where it can be more effective. Building on the allocation objective in Eq. (4), the optimization objective for the RL problem, i.e., the allocation reward, can be formulated as follows:
@@
We detail this RL problem under a Markov decision process framework, where the main components are:
1) State. We leverage all available information to represent the observations of each block’s prompts. Specifically, the state is defined as a concatenation of three elements: block-wise idleness scores, prompt distribution, and one-hot encoding of the pruned prompt’s location. Formally, the state is defined as $s\!=\!\{\sum_{d_{k}\!=\!i}\widehat{\mathcal{I}}_{k}\}_{i\!=\!1}^{L}\!\bigcup\!\mathcal{D}^{-}\!\bigcup\!\text{{OneHot}}(k^{\ast})$. After PR and VPT, the state transitions to $s^{\prime}$.
2) Action. The action determines which block the pruned prompt will be allocated to, represented as $a\in[L]$.
3) Reward. Although the allocation reward is defined in Eq. (7), calculating the intermediate loss of $f_{\mathbf{P},\mathcal{D}^{-}}$ is computationally inefficient. To address this, we reformulate the reward by incorporating the idleness score as:

### v2

# PRO-VPT: Distribution-Adaptive Visual Prompt Tuning via Prompt Relocation

@@ Abstract
Visual prompt tuning (VPT), i.e., fine-tuning some lightweight prompt tokens, provides an efficient and effective approach for adapting pre-trained models to various downstream tasks. However, most prior art indiscriminately uses a fixed prompt distribution across different tasks, neglecting the importance of each block varying depending on the task. In this paper, we introduce adaptive distribution optimization (ADO) by tackling two key questions: ① How to appropriately and formally define ADO, and ② How to design an adaptive distribution strategy guided by this definition? Through empirical analysis, we first confirm that properly adjusting the distribution significantly improves VPT performance, and further uncover a key insight that a nested relationship exists between ADO and VPT. Based on these findings, we propose a new VPT framework, termed PRO-VPT (iterative Prompt RelOcation-based VPT), which adaptively adjusts the distribution built upon a nested optimization formulation. Specifically, we develop a prompt relocation strategy derived from this formulation, comprising two steps: pruning idle prompts from prompt-saturated blocks, followed by allocating these prompts to the most prompt-needed blocks. By iteratively performing prompt relocation and VPT, our proposal can adaptively learn the optimal prompt distribution in a nested optimization-based manner, thereby unlocking the full potential of VPT. Extensive experiments demonstrate that our proposal significantly outperforms advanced VPT methods, e.g., PRO-VPT surpasses VPT by 1.6 pp and 2.0 pp average accuracy, leading prompt-based methods to state-of-the-art performance on VTAB-1k and FGVC benchmarks. The code is available at https://github.com/ckshang/PRO-VPT.
@@ 1 Introduction
Fine-tuning pre-trained vision models (PVMs) [26] has proven remarkably effective in adapting to a variety of downstream tasks [65, 14, 30]. However, the computational and storage costs associated with full fine-tuning are prohibitively high, particularly as model sizes continue to grow [40, 57, 1]. To overcome these challenges, parameter-efficient fine-tuning (PEFT) methods [61, 62, 20, 2], such as visual prompt tuning (VPT) [22], have emerged as more promising alternatives, garnering significant attention.
@@
Instead of updating all model parameters directly, VPT inserts lightweight learnable prompt tokens into the input space of Transformer blocks, with either a shallow distribution (inserted only in the first block) or a deep distribution (uniformly inserted across all blocks), while keeping the PVM frozen. This approach significantly reduces the number of parameters that need to be fine-tuned, thereby lowering the computational and memory overhead [14]. Follow-up research further enhances this pipeline by improving aspects such as prompting architecture [13, 50, 52], prompt initialization [58], and prompt propagation structure [71, 63, 5].
@@
Despite these advancements, most of the prior art still relies on a preset fixed prompt distribution, either shallow or deep, across various downstream tasks, as illustrated in Fig. 1. However, recent findings reveal that the importance of each pre-trained block varies significantly across different tasks [49, 43, 12, 28, 60], underscoring the need to dynamically tailor the fine-tuning intensity for each block to improve task performance [29, 53, 11, 45, 25]. This implies that indiscriminately using a fixed prompt distribution may not fully exploit the potential of VPT, motivating us to adaptively adjust the distribution to better accommodate each task.
@@
In this paper, we naturally consider prompt distribution as one of the optimization objectives, which we refer to as adaptive distribution optimization (ADO). As the focus pivots towards the ADO problem, new questions arise. First, since the underlying nature of distribution optimization is still unexplored, the definition of ADO remains unclear. For instance, defining it as a one-shot process or an iterative one leads to a critical consideration, namely whether the effectiveness of distribution adjustments is influenced by ongoing prompt updates. This leads us to our first key question: ① How can we appropriately and formally define ADO?3.2) Second, once the definition of ADO is established, a more essential question emerges: ② How can we achieve adaptive distribution derived from this definition?3.3)
@@
To address these questions, we first empirically analyze the ADO problem. Our findings confirm the significance of ADO and further reveal two key insights: the effectiveness of distribution adjustments is indeed impacted by prompt updates, and a nested relationship exists between ADO and VPT. Based on these insights, we propose PRO-VPT (iterative Prompt RelOcation-based VPT), a novel iterative framework derived from a nested optimization definition. Specifically, building on this definition, we develop a prompt relocation (PR) strategy for ADO, which comprises two optimization steps: pruning and allocating. The pruning step first identifies and releases idle prompts from prompt-saturated blocks, while the subsequent allocation step determines the most prompt-needed blocks for redistributing these prompts. By iteratively performing PR and VPT, PRO-VPT can adaptively learn the optimal prompt distribution for each task in a nested optimization manner, thereby unlocking the full potential of VPT. Our main contributions are fourfold.
@@
  • Derived from the definition, we present a PR strategy for ADO, which introduces an idleness score to prune underutilized prompts from prompt-saturated blocks and then allocates them to prompt-needed blocks using reinforcement learning (RL).
  • Extensive experiments show that PRO-VPT exhibits superior performance, e.g., it attains state-of-the-art average accuracies of 78.0% on VTAB-1k and 91.7% on FGVC, with 1.6 pp and 2.0 pp advantages over the VPT baseline.
@@ 2 Related Work
Visual Prompt Tuning. VPT [22], one of the leading techniques in PEFT, draws inspiration from language-domain prompting works [37, 32, 38, 19] to fine-tune a small portion of prompts while keeping the pre-trained model frozen. Subsequent research has improved VPT in three main aspects: (1) Extending prompts to other Transformer components (e.g., key and value matrices or the patch embedding layer) to better suit the Transformer architecture [13, 50, 52]; (2) Initializing prompts with prototypes of image tokens to accelerate convergence [58]; (3) Refining the propagation structures of prompts to strengthen cross-layer interactions [71, 63, 5]. However, most existing methods depend on a fixed prompt distribution for all tasks, resulting in less effective prompting performance. In contrast, our approach adaptively adjusts the prompt distribution for each task, enabling prompt-based methods to achieve state-of-the-art performance.
Layer-wise Fine-Tuning. Recent research has highlighted the importance of treating each pre-trained block differently during downstream fine-tuning, emphasizing the need for layer-wise adaptive strategies to calibrate the fine-tuning intensity for each block [54, 31, 12, 43, 25]. For instance, [29] first revealed that some pre-trained blocks may already be near-optimal for specific tasks, indicating that fine-tuning demands of each block vary significantly for different tasks. This work then proposes freezing these near-optimal blocks and selectively fine-tuning the remaining ones to improve performance. Similarly, [49, 45] proposes to adaptively assign different learning rates to each block, tailoring the fine-tuning intensity at a layer level to achieve more fine-grained layer-wise adaptation. Inspired by these findings, we hypothesize that the prompting intensity for each block, i.e., the overall distribution of prompts, may also play a crucial role in task adaptation, leading us to focus on adaptively adjusting the prompt distribution within VPT.
@@ 3 Methodology
3.1 Notation
Visual prompt tuning (VPT) aims to learn lightweight tunable prompts $\mathbf{P}$ that are inserted into the vision Transformer (ViT) $f(\cdot)$ with a specific distribution $\mathcal{D}$. In particular, for a set of $N$ prompts $\mathbf{P}\!=\!\{\mathbf{p}_{k}\}^{N}_{k=1}$ and their distributed blocks $\mathcal{D}\!=\!\{d_{k}\}^{N}_{k=1}$ where $d_{k}\!\in\!\{0,1,\ldots,L\}$ ($d_{k}\!=\!0$ for undistributed), the prompted ViT model is denoted as $f_{\mathbf{P},\mathcal{D}}(\cdot)$. Given a training set $\mathcal{T}_{tr}$ of input-label pairs $(\mathbf{x},y)$, the model is then trained by minimizing the task loss $\mathcal{L}(f_{\mathbf{P},\mathcal{D}}(\mathbf{x}),y)$. For convenience in later sections, we define the loss difference as $\Delta\mathcal{L}(f_{1},f_{2})\!=\!\mathcal{L}(f_{1}(\mathbf{x}),y)\!-\!\mathcal{L}(f_{2}(\mathbf{x}),y)$.
@@ 3.2 Analysis and Formulation of ADO
We first explore the underlying nature of ADO and its formal definition, in response to question ①.
Intuitively, ADO can be defined in two distinct ways: one-shot (independently adjusting the distribution) or iterative (integrating with prompt tuning through iteratively adjusting the distribution). These definitions are distinct in whether the distribution is adjusted concurrently with prompt tuning. Specifically, an iterative process is more appropriate when distribution adjustments are affected by ongoing changes in prompts, as independent optimization may lead to suboptimal results by failing to account for these dynamics. Otherwise, one-shot adjustments may be sufficient.
@@
To investigate which definition better matches the underlying nature, we experiment to examine how distribution adjustments behave when applied to prompts from different epochs. We use two intuitive adjustment strategies: incrementally adding to a small initial prompt set and repositioning existing prompts. As shown in Fig. 2, adding or repositioning prompts to specific blocks can effectively improve performance; however, the effective adjustments differ significantly when prompts are updated. Comprehensive results with detailed analysis are provided in Appendix C. In conclusion, we summarize our key findings as follows:
Finding 1. Properly adjusting the prompt distribution significantly enhances performance, confirming the necessity to reconsider its optimization.
Finding 2. Distribution adjustments are sensitive not only to different blocks but also to the updated prompts themselves, highlighting the need for an iterative process that continuously adjusts the distribution during training.
Finding 3. The effectiveness of distribution adjustments can only be properly evaluated after prompt tuning, indicating that the prompt tuning process should be nested within the distribution optimization process (see Appendix C).
Building on these insights, we propose integrating ADO with VPT into an iterative optimization framework. Formally, we present the definition of this ADO-VPT co-design framework through the lens of nested optimization, where the objectives of ADO and VPT are expressed as:
$\mathcal{D}^{\ast}$ $=\mathop{\arg\min}\limits_{\mathcal{D}}\penalty 10000\ \mathbb{E}_{(\mathbf{x},y)\in\mathcal{T}_{tr}}\big[\mathcal{L}(f_{\mathbf{P}^{\ast},\mathcal{D}}(\mathbf{x}),y)\big],$ (1)
$\mathbf{P}^{\ast}$ $=\mathop{\arg\min}\limits_{\mathbf{P}}\penalty 10000\ \mathbb{E}_{(\mathbf{x},y)\in\mathcal{T}_{tr}}\big[\mathcal{L}(f_{\mathbf{P},\mathcal{D}^{\ast}}(\mathbf{x}),y)\big].$ (2)
This formulation clearly illustrates the intertwined relationship between visual prompts $\mathbf{P}$ and their distribution $\mathcal{D}$, with its framework presented in Fig. iii. In this framework, prompt tuning is integrated into the ADO process, forming a closed-loop workflow of ‘distribution adjustment $\rightarrow$ prompt tuning $\rightarrow$ adjustment evaluation $\rightarrow$ new adjustment’, thereby achieving a co-design process aligning with the observed iterative and nested nature.
Remark. This nested optimization framework offers a universal approach capable of incorporating various distribution adjustment strategies, e.g., adding or repositioning. However, we empirically observe that adding-based adjustments often lead to unstable performance (see Appendix D). Therefore, in the following section, we focus on developing a prompt relocation (PR) strategy based on this formulation to effectively tackle the ADO problem.
@@ 3.3 Iterative Prompt Relocation-based VPT
Following the relocation objective defined in Eq. (1), we introduce the PR strategy derived from this definition and provide a detailed overview of the PRO-VPT framework, in response to question ②.
Building on the aforementioned findings, the discrete PR objective is further complicated by the non-stationary nature of effective relocation (Finding 2) and the necessity to predict the effectiveness of relocation operations (Finding 3). Nonetheless, these three properties make a RL framework [33, 24] a natural and well-suited approach for optimizing the PR objective. However, a naive RL implementation needs to model all possible relocation operations as actions, resulting in $L^{2}$ possible arrangements (e.g., 144 actions for ViT-B). Such a large action space necessitates numerous iterations to achieve convergence (cf. § 4.3).
To address this, we propose decomposing the PR process into two sequential steps: pruning and allocating. The pruning step first identifies and releases idle prompts for relocation, while the allocating step then employs RL to determine optimal blocks for these freed prompts. By focusing solely on allocatable blocks instead of all block arrangements, this approach effectively reduces the action space from $L^{2}$ to $L$. More importantly, this approach has an intuitive explanation: idle prompts typically exist within the prompt-saturated blocks, thus relocating them to prompt-needed blocks may enhance the effectiveness of prompt distribution. Decomposed from the relocation objective in Eq. (1), the pruning and allocation steps [^1] Here, for simplicity and efficiency, both the pruning and allocation steps are formalized to operate on a single prompt per iteration, as further elaborated in Appendix E. can be formalized as:
$k^{\ast}=\$$\mathop{\arg\max}\limits_{k}\penalty 10000\ \mathbb{E}_{(\mathbf{x},y)\in\mathcal{T}_{tr}}\big[\Delta\mathcal{L}(f_{\mathbf{P},\mathcal{D}},f_{\mathbf{P},\mathcal{D}\mid d_{k}=0})\big],$ (3)
$a^{\ast}=\$$\mathop{\arg\max}\limits_{a}\penalty 10000\ \mathbb{E}_{(\mathbf{x},y)\in\mathcal{T}_{tr}}\big[\Delta\mathcal{L}(f_{\mathbf{P},\mathcal{D}^{-}},f_{\mathbf{P}^{\prime},\mathcal{D}\mid d_{k^{\!\ast\!}}=a})\big],$ (4)
where $\mathcal{D}^{-}\!=\!\{d_{k}\!\in\!\mathcal{D}\!\mid\!d_{k^{\!\ast\!}}\!=\!0\}$ and $\mathcal{D}^{+}\!=\!\{d_{k}\!\in\!\mathcal{D}\!\mid\!d_{k^{\!\ast\!}}\!=\!a^{\ast}\}$ denote the prompt distributions after pruning and allocation with $a\!\in\![L]$ as the allocation block, and the allocation step uses the tuned prompts $\mathbf{P}^{\prime}$ for evaluation following Finding 3. These two equations are specifically designed to identify the $k^{*}$-th prompt suitable for pruning and to determine the optimal block $a^{*}$ for allocating.
Following this, we utilize Eqs. (3) and (4) to approach the pruning and allocation stages, respectively, with the overall PRO-VPT framework illustrated in Fig. 3 and Algo. 1:
Idleness Score-based Pruning. To measure whether a prompt is idle for its current block and suitable for relocation, we take Eq. (3) as the idleness score:
@@
Intuitive Interpretation. If $\mathcal{I}_{k}\!=\!\Delta\mathcal{L}\!>\!0$, pruning the prompt $\mathbf{p}_{k}$ from its current block can enhance performance. This indicates that $\mathbf{p}_{k}$ is negative for its current block, which tends to be prompt-saturated and therefore suitable for its relocation to a more prompt-needed block. Conversely, $\mathcal{I}_{k}\!\leq\!0\penalty 10000\ \forall k$ suggests that no blocks are saturated and thus the prompt distribution is near-optimal.
Pruning Implementation. Since computing $\mathcal{I}_{k}$ for each individual prompt $\mathbf{p}_{k}$ is significantly inefficient, we approximate $\mathcal{I}_{k}$ using its first-order Taylor expansion (see Appendix F for detailed derivation):
$\widehat{\mathcal{I}}_{k}\approx\mathbf{g}_{k}^{T}\mathbf{p}_{k},$ (6)
@@
To this end, by calculating the estimated scores as in Eq. (6), we can effectively identify and prune the idle prompt $\mathbf{p}_{k^{\!\ast\!}}$ with the maximum $\widehat{\mathcal{I}}_{k^{\!\ast\!}}$ from the prompt-saturated block (when $\max\widehat{\mathcal{I}}_{k}\!>\!0$), thereby fulfilling Eq. (3) and forming the pruned distribution $\mathcal{D}^{-}$.
RL-based Allocation. After pruning an idle prompt, the following key step is to leverage RL to determine its optimal allocation to the prompt-needed block. Specifically, we take Eq. (4) as the optimization objective for the RL problem, i.e., the allocation reward, as follows:
@@
Intuitive Interpretation. The optimal allocation block $a^{\ast}$, which achieves the maximum $r^{\ast}\!=\!\max\Delta\mathcal{L}$, is identified as the block where assigning a prompt leads to the highest performance improvement, implying that this block is the most prompt-needed.
Allocation Implementation. We detail the RL-based allocation under a Markov decision process framework, where the main components are:
(1) State. We leverage all available information to represent the state of the current prompt distribution. Specifically, the state incorporates the block-wise idleness scores, the current distribution, and the pruned prompt’s location, formally defined as $s\!=\!\{\sum_{d_{k}\!=\!i}\widehat{\mathcal{I}}_{k}\}_{i\!=\!1}^{L}\!\bigcup\!\mathcal{D}^{-}\!\bigcup\!\text{{OneHot}}(k^{\ast})$. After PR and VPT, the state transitions to $s^{\prime}$.
(2) Action. The action determines which block the pruned prompt will be allocated to, represented as $a\in[L]$.
(3) Reward. Although the allocation reward $r$ is defined in Eq. (7), calculating the intermediate loss of $f_{\mathbf{P},\mathcal{D}^{-}}$ is computationally inefficient. To address this, we approximate $r$ by incorporating the idleness score as: